# Aligning Perception, Reasoning, Modeling and Interaction: A Survey on Physical AI

### v3

@@ Abstract
The rapid advancement of embodied intelligence and world models has intensified efforts to integrate physical laws into AI systems, yet physical perception and symbolic physics reasoning have developed along separate trajectories without a unified bridging framework. This work provides a comprehensive overview of physical AI, establishing clear distinctions between theoretical physics reasoning and applied physical understanding while systematically examining how physics-grounded methods enhance AI’s real-world comprehension across structured symbolic reasoning, embodied systems, and generative models. Through rigorous analysis of recent advances, we advocate for intelligent systems that ground learning in both physical principles and embodied reasoning processes, transcending pattern recognition toward genuine understanding of physical laws. Our synthesis envisions next-generation world models capable of explaining physical phenomena and predicting future states, advancing safe, generalizable, and interpretable AI systems. We maintain a continuously updated resource at https://github.com/AI4Phys/Awesome-AI-for-Physics.
@@ A Introduction
Teaching artificial intelligence to understand our physical world represents one of the most fundamental challenges in modern AI research [1, 2]. While humans intuitively predict object dynamics and grasp complex physical interactions from early childhood, current AI models struggle with basic physical reasoning that young children master effortlessly [3, 4]. This capability gap becomes increasingly critical as AI systems are deployed in real-world scenarios ranging from autonomous vehicles navigating dynamic environments to robotic systems manipulating objects in unfamiliar settings.
@@
Recent large-scale benchmarks have emerged to quantify this fundamental limitation. Evaluation frameworks such as SeePhys [5] for symbolic reasoning, PHYRE [6] for intuitive physical perception, and PhyBlock [7] for realistic dynamics prediction consistently demonstrate that current models lack structured understanding of physical laws governing our universe. Remarkably, a vision model trained on millions of images may achieve superhuman object recognition accuracy yet fail to predict elementary outcomes such as a bouncing ball’s trajectory or the stability of stacked blocks [8, 9]. This paradox reveals a fundamental issue: contemporary AI models primarily learn statistical correlations from data rather than developing causal understanding rooted in physical principles.
@@
Figure 1: Overview of four physical understanding capabilities of current AI systems. Physical Perception denotes model’s ability to recognize the properties of objects; Physics Reasoning refers to model’s capability to understand symbolic systems and physical laws; World Modeling represents its capacity for simulation and prediction; and Embodied Interaction characterizes model’s ability to interact with the physical environment.
To address this challenge, the research community has pursued physics-aware reasoning models across multiple complementary directions. From an architectural perspective, Graph Neural Networks (GNNs) [10, 11, 12] have emerged as particularly promising, leveraging their inherent capacity to model relational structures and capture pairwise physical interactions between objects [4, 13, 14]. These graph-based approaches naturally encode spatial relationships and interaction dynamics, making them well-suited for physical reasoning tasks. Concurrently, diffusion-based models [15, 16] have shown remarkable potential in modeling complex physical processes through learned denoising procedures, while transformer architectures [17, 18] increasingly incorporate physics-aware attention mechanisms and constraints to enhance physical understanding.
From a training methodology perspective, researchers have developed physics-guided loss functions that explicitly incorporate physical constraints [19], curriculum learning strategies that progressively introduce physical complexity [20], and reinforcement learning frameworks with physics-informed reward structures [21, 22]. These training paradigms aim to instill physical intuition during the learning process rather than relying solely on pattern recognition.
For inference enhancement, recent approaches explore the integration of explicit physical rules and symbolic reasoning [23, 24, 25], as well as differentiable physics engines that enable end-to-end optimization [26, 27]. These methods bridge the gap between learned representations and established physical principles.
A particularly promising emerging pathway leverages large language models (LLMs) and multimodal large language models (MLLMs). Recent breakthroughs have demonstrated their impressive capability in learning from massive datasets and performing sophisticated end-to-end reasoning across multiple modalities [28, 29, 30, 31, 32, 33]. Notably, works employing multi-agent systems and tool integration have achieved performance comparable to top human competitors in physics competitions [qiu2024physics, 34], suggesting the potential for more general physical understanding.
While existing surveys have provided valuable insights into specific sub-domains [35, 36, 37, 38], the rapidly evolving field requires a comprehensive review that systematically tracks the evolution of AI models across the full spectrum of physics-oriented challenges.
Contributions. We present a comprehensive three-tier taxonomy that systematically captures reasoning capabilities and downstream applications in the current landscape of physical AI systems. We categorize contemporary research efforts into four fundamental capabilities which are shown in Figure 1: physical perception (drawing conclusions from observations of physical phenomena), physics reasoning (utilizing symbolic representations and mathematical methods to solve theoretical physics problems), world modeling (modeling and predicting dynamics in virtual environments that mirror our physical world), and embodied interaction (navigating and manipulating real objects in physical environments).
For each capability domain, we organize tasks into hierarchical levels and provide detailed analysis of representative implementation methodologies for each application scenario. Drawing from over 300 papers, our comprehensive coverage spans from research-oriented problems in advanced physics (e.g., quantum mechanics, statistical physics) to application-oriented tasks including object recognition, spatial perception, video generation, robotic control, and autonomous driving systems.
@@
Survey Structure. Section B establishes our scope, presents the proposed taxonomy, and reviews prerequisite knowledge including model architectures and training algorithms. Sections D through F systematically examine each reasoning capability with corresponding tasks and methodological approaches. Section G outlines promising future directions toward more capable physical AI systems. Section H synthesizes our contributions and findings.
@@
B Preliminaries
B.1 Scope
We focus on how frontier AI systems tackle physics and physical reasoning problems in the era of deep learning. For physics reasoning, we specifically focus on how AI systems leverage the laws of physics to solve physics problems in end to end manner, which is as opposed to how physics inspired AI development. As we aim to provide a coherent discussion on a central thesis, we do not discuss how physics inspire machine learning methods (such as the Boltzmann Machine and Hopfield Network) or how machine learning at large facilitates physics research, which is a vast and diversified field we leave for future work in AI for Science.
B.2 Definition
Large Language Models. Large Language Models (LLMs) [39, 33, 40, 41, 42] are neural networks trained on vast text corpora to predict the next token in a sequence, enabling them to generate coherent text and perform various language tasks. LLMs employ transformer architectures [43] to capture long-range dependencies through self-attention mechanisms: $\text{Attention}(Q,K,V)=\text{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V$. Mathematically, LLMs learn:
$p_{\theta}(x)=\prod_{i=1}^{n}p_{\theta}(x_{i}|x_{<i})$ (1)
where $x_{i}\in\mathcal{V}$ represents tokens from vocabulary $\mathcal{V}$. Multimodal Large Language Models (MLLMs) [33, 31, 28, 29] extend LLMs to process and reason about multiple modalities, e.g. images, audio and video, within a unified framework. MLLMs achieve this through:
$p_{\theta}(x,v)=p_{\theta}(x|f_{\phi}(v))\cdot p(v)$ (2)
where $f_{\phi}:\mathcal{I}\rightarrow\mathbb{R}^{d}$ encodes visual input $v$ into the language model’s representation space, enabling vision-language understanding through shared embeddings [44]. Both LLMs and MLLMs are optimized by minimizing the negative log-likelihood $\mathcal{L}=-\sum_{i}\log P(x_{i}|x_{<i})$.
Visual Generation Models. Image and video generation models are deep learning systems that learn to synthesize realistic visual content by modeling the complex distribution of natural images and videos. These models are mostly trained on large scale visual datasets, enabling them to generate content from noise or conditional inputs like text descriptions, sketches, or partial observations. A unified formulation for both:
$p_{\theta}(x_{0:T}|c)=\prod_{t=0}^{T}p_{\theta}(x_{t}|x_{<t},c),$ (3)
where $x_{t}\in\mathbb{R}^{H\times W\times C}$ represents the visual frame at time $t$, $T=0$ for single image generation and $T>0$ for video sequences, and $c$ encodes conditioning signals (text prompts, class labels, initial frames, or latent codes). Modern implementations typically use diffusion models, GANs, or autoregressive transformers to approximate this distribution.
3D Reconstruction Models. 3D reconstruction models recover three-dimensional structure and appearance from 2D observations such as images or videos. These models solve the inverse rendering problem by inferring geometry (shape, depth, surface normals) and appearance properties (color, texture, materials) from partial visual data, often leveraging multi-view consistency, learned shape priors, or physics-based constraints. Formally: $\mathcal{M}_{\theta}:\{\mathcal{I}_{i},\mathcal{K}_{i}\}_{i=1}^{N}\rightarrow(\mathcal{G},\mathcal{A})$ where $\mathcal{I}_{i}\in\mathbb{R}^{H\times W\times 3}$ are input images with camera parameters $\mathcal{K}_{i}\in SE(3)$, outputting geometry $\mathcal{G}$ (represented as meshes, point clouds, voxels, or implicit fields $f:\mathbb{R}^{3}\rightarrow\mathbb{R}$) and appearance $\mathcal{A}$ (texture maps, radiance fields, or material properties).
Vision Language Action Models. Vision-Language-Action (VLA) models are embodied AI systems that directly map visual observations and language instructions to continuous control actions for physical manipulation tasks. These models enable robots to perform complex tasks by understanding natural language commands and visual context, outputting precise motor controls. Formally: $a_{t}=\pi_{\theta}(I_{t},l,h_{t})$ where $a_{t}\in\mathbb{R}^{d_{a}}$ represents continuous actions (e.g., end-effector position, joint angles), $I_{t}$ is the visual input (RGB/depth images), $l$ is the language instruction, and $h_{t}$ captures temporal context or proprioceptive state.
Vision Language Navigation Models. Vision-Language Navigation (VLN) models enable agents to navigate through environments by following ninstructions while processing visual observations. VLN typically involves discrete action selection from a predefined set of navigation primitives. Formally: $a_{t}=\arg\max_{a\in\mathcal{A}_{nav}}\pi_{\theta}(o_{t},l_{inst},m_{t})$ where $\mathcal{A}_{nav}=\{\text{forward, left, right, stop}\}$ represents discrete navigation actions, $o_{t}$ is the current observation (panoramic or egocentric view), $l_{inst}$ is the navigation instruction, and $m_{t}$ encodes spatial memory or topological map representations.
Mathematical Reasoning. Mathematical reasoning has long served as the foundation of scientific inquiry. In the context of artificial intelligence, advances in automated mathematical reasoning have begun to influence the study of physical reasoning, enabling models not only to manipulate symbolic structures but also to ground them in physical laws and quantitative relations [45, 46, 47, 48]. Mathematical understanding requires LLMs to parse and contextualize diverse structures ranging from arithmetic operations to geometric relationships while integrating textual and visual representations. Early explorations [49] relied on manually crafted rules and pattern-matching heuristics, which evolved into semantic parsing approaches [50, 51] that translated natural-language problems into structured logical forms. With the advent of large-scale pretrained models such as ChatGPT [40], the field underwent a paradigm shift [52] by scaling laws [53, 54] focusing on training LLMs with large-scale heterogeneous datasets [55, 56] to enable better reasoning generalization.
The Rise of CoT-based thinking models. Chain-of-Thought (CoT) [57] has emerged as a pivotal technique by encouraging models to produce intermediate reasoning steps prior to delivering final responses. A substantial body of research has demonstrated that explicit generation of such step-by-step reasoning significantly enhances model performance, particularly in domains requiring logical inference. Advances exemplified by OpenAI o1 [30] and DeepSeek R1 [39], have introduced extended CoT reasoning through test-time scaling, enabling more comprehensive, structured, and iterative reasoning process.
Supervised Fine-tuning and Distillation. Supervised Fine-tuning (SFT) adapts pre-trained models to specific tasks by optimizing on labeled input-output pairs (Equation 4), and when combined with knowledge distillation, creates a powerful framework where strong teacher models generate high-quality training data to improve smaller student models’ reasoning capabilities. As demonstrated by previous work [58, 59, 60, 61], this SFT-based distillation of reasoning traces from teacher to student models yields substantial gains in mathematical reasoning proficiency by transferring not only final answers but also nuanced reasoning pathways and strategic problem-solving behaviors. Recent efforts [62, 63] have further enhanced this approach by incorporating error-corrected data into the SFT pipeline, exposing models to both flawed reasoning paths and their corrections to foster robustness and self-diagnostic capabilities.
$\mathcal{L}_{\text{SFT}}=-\mathbb{E}_{(x,y)\sim D}[\text{log}(\pi_{\theta}(y|x))].$ (4)
Reinforcement Learning. Notably, DeepSeek-R1 [39] integrates reinforcement learning into the reasoning process with a structured output format, which explicitly separates internal reasoning traces from final answer output (e.g., The optimization method is given in Equation 5), which greatly enhanced the interpretability and reliability of model outputs. Complementing such architectural innovations, pioneering work also relied on rule-based reward shaping to guide reasoning trajectories; yet these methods often suffered from LLM hallucination or logically inconsistent intermediate steps. In response, Process Reward Models (PRMs) [64, 65] have been developed to provide fine-grained supervision over each reasoning step, thereby improving faithfulness and coherence across the entire chain of thought.
$\begin{split}\mathcal{L}_{\text{RL}}=\mathbb{E}\Big[&\min\left(r_{t}(\theta)A_{t},\text{clip}(r_{t}(\theta),1-\varepsilon,1+\varepsilon)A_{t}\right)\\ &-\beta D_{\text{KL}}[\pi_{\theta}\|\pi_{\text{old}}]\Big]\end{split}$ (5)
External Verification. Another research direction is to overcome inherent capacity constraints of standalone models, recent work [66, 67, 68, 69, 70, 71] explore the use of external tools that provide signals for factual or logical soundness. These include retrieval of external knowledge bases [68], image pre-processing operations such as zooming and cropping in multimodal settings [69], and the use of auxiliary visual guides [70, 71]. Such hybrid architectures extend the functional boundaries of LLMs by integrating complementary systems, mimicking human-like resourcefulness in problem solving. Collectively, these advancements reflect a paradigm shift from static inference toward dynamic, reflective, and interactive reasoning. From long CoT and PRM-guided training to reinforcement learning using various self-evolutionary mechanisms with verifications signals, the field is rapidly progressing toward increasingly autonomous and reliable AI reasoners.
@@ B.2 Definition
B.3 Taxonomy
To systematically understand how AI systems engage with physics, we propose a hierarchical taxonomy that reflects the natural progression from abstract reasoning to embodied action. Our framework organizes physics-aware AI capabilities into four interconnected domains that mirror human cognitive development in physical understanding. Physical Perception represents the ability to intuitively extract physical property from sensory data through increasingly sophisticated layers, including object recognition, spatial relationships and intrinsic. In the the inverse direction, Physics Reasoning encompasses symbolic manipulation and theoretical problem-solving, where AI systems leverage mathematical representations to solve physics problems ranging from textbook exercises to research-level challenges. These two complementary pathways converge in World Modeling, where AI systems integrate symbolic knowledge with perceptual understanding to construct predictive models of physical environments, enabling capabilities from video generation to reconstruction. Finally, Embodied Interaction grounds these cognitive capabilities in real-world action, where robots, autonomous vehicles, and navigation agents must reconcile theoretical understanding with the irreversible consequences of physical manipulation. Figure 3 shows our taxonomy. This taxonomy not only delineates current research boundaries but also reveals critical gaps where symbolic reasoning and embodied experience remain disconnected, highlighting the path toward genuinely physics-aware artificial intelligence.
C Physical Perception
Understanding the physical world begins with perception. Before AI systems can reason about abstract physical laws or manipulate objects in complex environments, they must first acquire the fundamental capability to perceive, understand, and infer physical properties from sensory inputs. We term this capability Physical Perception, which represents the foundational layer of physical intelligence, where raw sensory data from cameras, sensors, and other modalities are transformed into structured representations of the physical world. In this section, we organize existing investigations into five hierarchical categories: Object Recognition, Spatial Perception, Intrinsic Property, Dynamic Estimation, and Causal & Counterfactual Inference, ordered by increasing demands of cognitive load.

### v4

@@ Abstract
The convergence of embodied intelligence and world models has catalyzed growing interest in integrating physical laws into AI systems. While prior surveys have examined world models and embodied intelligence separately, we focus on the progression that connects these capabilities as a unified developmental pathway from passive observation to active physical comprehension. This survey provides a systematic framework revealing how physical AI advances through four interconnected stages: perception transforms sensory data into structured physical representations, reasoning derives explanations from observed phenomena, modeling enables predictive simulation grounded in physical principles, and embodied interaction closes the loop through physical manipulation and environmental feedback. Each stage enables and enhances the next: perceptual grounding supports causal reasoning, reasoning unlocks predictive capabilities, and robust models drive genuine physical interaction. Through analysis of developments spanning architectural innovations, training methodologies, causal inference, and embodied systems, we synthesize how physical understanding emerges through cumulative integration across this progression. Our framework reveals the evolution from isolated, task-specific solutions toward integrated architectures that advance from pattern recognition toward causal reasoning and counterfactual prediction. This perspective provides foundations for next-generation physical AI systems with direct implications for safe, generalizable, and interpretable deployment across robotics, scientific discovery, and autonomous systems. We maintain a continuously updated taxonomy repository at https://github.com/AI4Phys/Awesome-AI-for-Physics.
Index Terms:
Physical AI System, Physical Perception, Physics Reasoning, World Modeling, Embodied Interaction.
@@ A Introduction
Teaching artificial intelligence to understand our physical world represents one of the most fundamental challenges in modern AI research [World_models2018, AIMeetPhy_survey_2024]. While humans naturally grasp complex physical interactions from early childhood, frontier models struggle with basic physical reasoning that young children master effortlessly [NewtonianIU, Battaglia2016InteractionNF]. This capability gap becomes increasingly critical as AI systems are deployed in real-world scenarios ranging from self-driving vehicles to humanoid robots. To quantify this limitation, evaluation frameworks such as SeePhys [seephys] for symbolic reasoning, PHYRE [I-PHYRE] for intuitive physical perception, and PhyBlock [Ma2025PhyBlockAP] for realistic dynamics prediction have emerged, consistently demonstrating that current models lack structured understanding of physical laws. This gap manifests dramatically in performance disparities. A vision model trained on millions of images may achieve superhuman image classification accuracy yet fail to predict elementary outcomes such as a bouncing ball’s trajectory [LLMPhy_2025, ComPhy]. This paradox reveals that contemporary AI models learn statistical correlations from massive data rather than developing causal knowledge rooted in physical principles.
@@
Figure 1: Overview of four physical understanding capabilities of current AI systems. Physical Perception denotes the model’s ability to recognize the properties of objects. Physics Reasoning refers to the model’s capability to understand symbolic systems and physical laws. World Modeling represents its capacity for simulation and prediction. Embodied Interaction characterizes the model’s ability to interact with the physical environment.
Addressing this fundamental challenge requires physics-aware reasoning models across multiple complementary directions. From an architectural perspective, Graph Neural Networks (GNNs) [gcn, gat, graphsage] leverage their inherent capacity to model relational structures and capture pairwise physical interactions between objects [Battaglia2016InteractionNF, VIN, NPE], naturally encoding spatial relationships and interaction dynamics. Diffusion-based models [MotionCraft, VideoREPA] have demonstrated substantial potential in modeling complex physical processes through learned denoising procedures, while transformer architectures [Kantamneni2024HowDTA, Han2025APMA] increasingly incorporate physics-aware attention mechanisms and constraints. Beyond architecture, training methodology plays a crucial role. Researchers have developed physics-guided loss functions that explicitly incorporate physical constraints [elhasadi2019_semisupervised_sr_drag], curriculum learning strategies that progressively introduce physical complexity [chen2025mint], and reinforcement learning frameworks with physics-informed reward structures [AlphaDrive, AutoDrive, liu2025reinforcementlearningmeetslarge]. These training paradigms aim to instill physical intuition during learning rather than relying solely on pattern recognition. At test-time inference, approaches explicitly integrate physical laws and symbolic reasoning into causal modeling [causalmodeling, causalthreads, PhysORD2024], while differentiable physics engines enable end-to-end optimization [differentiable-stokes-flow2020, Scalable-differentiable-physic2020], thereby bridging the gap between learned representations and established physical principles.
Among these directions, large language models (LLMs) and multimodal large language models (MLLMs) offer a particularly promising pathway forward. These models have demonstrated substantial capability in learning from massive datasets and performing sophisticated end-to-end reasoning across multiple modalities [gpt4o, llava, o1, Qwen2.5-VL, Claude-3.7-Sonnet, Gemini-pro]. Multi-agent systems with tool integration have achieved performance comparable to top human competitors in physics competitions [Liang2025MultimodalRF], suggesting potential for more general physical understanding. Generative systems [GAIA-1, DriveDreamer] have achieved controllable, photorealistic synthesis of physical scenarios, while vision-language-action models [KimEtAl2024_OpenVLA, BlackEtAl2024_pi0] integrate natural language instructions with continuous physical manipulation, bridging the gap from abstract reasoning to embodied interaction. These developments reflect a broader paradigm shift in physical AI from isolated, task-specific solutions toward integrated architectures (Figure 2) with synergies across domains, extending beyond recognizing input features toward generating and interacting with realistic physical scenarios.
Scope Comparison and Contributions. As summarized in Table I, existing surveys have examined individual dimensions of physical understanding in isolation, addressing perception [Yin2023ASO, Marjieh2023LargeLM, Sapkota2025ObjectDW, Guo2024PhyGraspGR, Sun2025ProbingPC], reasoning [Zhou2025FromPT, Li2025FromS1, Ravishankara2025TheAI, Duan2022ASO, Huang2022TowardsRI], modeling [Khan2025FoundationMD, Barman2025LargePM, Ding2024UnderstandingWO, Liu2025GenerativePA, Zhu2024IsSA, Kong20253DA4, Bergen2019MachineLF, Xie2025From2T, Chen2025ASO, Mai2024FromEM], and interaction [Liu2024AligningCS, Qi2024ShapeLLMU3, Sathyam2025FoundationMF, Feng2025EmbodiedAF, Long2025ASL, Wang2023ASO, Liang2025LargeME, Zheng2024ASO, Wang2025TowardEA, Duan2021ASOA, Xu2024ASOA, Han2023ASOA] as separate research areas without examining the synergistic connections between them. Our survey uniquely focuses on the evolutionary trajectory that unites these four capabilities into a coherent paradigm, analyzing how they interact and inform one another toward unified physical AI systems. We examine how deep learning systems leverage the laws of physics to solve physics problems in an end-to-end manner, as opposed to how physics principles inspire neural network architectures. We leave physics-inspired architectures such as Boltzmann Machines and Hopfield Networks, as well as the broader field of deep learning for physics research, for dedicated AI4Science surveys. To this end, we present a three-tier taxonomy that systematically organizes research across these four fundamental capabilities (Figure 1), with hierarchical task structures and detailed methodological analysis for each. Drawing from over 300 papers, our coverage spans advanced physics problems to applied tasks including object recognition, spatial perception, video generation, robotic control, and autonomous driving, revealing the integrated development pathway from passive observation to active physical comprehension.
@@
Survey Structure. Section B presents the proposed taxonomy. Sections D through F systematically examine each reasoning capability with corresponding tasks and methodological approaches. Section G outlines promising future directions toward more capable physical AI systems. Section H synthesizes our contributions and findings.
@@
TABLE I: Scope of existing surveys. They tend to examine only single dimensions of physical understanding abilities, without providing a comprehensive overview of physical AI evolution.
Capability Existing Surveys
Perception [Yin2023ASO, Marjieh2023LargeLM, Sapkota2025ObjectDW, Guo2024PhyGraspGR, Sun2025ProbingPC]
Reasoning [Zhou2025FromPT, Li2025FromS1, Ravishankara2025TheAI, Duan2022ASO, Huang2022TowardsRI]
Modeling [Khan2025FoundationMD, Barman2025LargePM, Ding2024UnderstandingWO, Liu2025GenerativePA, Zhu2024IsSA, Kong20253DA4, Bergen2019MachineLF, Xie2025From2T, Chen2025ASO, Mai2024FromEM]
Interaction [Liu2024AligningCS, Qi2024ShapeLLMU3, Sathyam2025FoundationMF, Feng2025EmbodiedAF, Long2025ASL, Wang2023ASO, Liang2025LargeME, Zheng2024ASO, Wang2025TowardEA, Duan2021ASOA, Xu2024ASOA, Han2023ASOA]
B Taxonomy
@@
To systematically understand how AI systems engage with physics, we propose a hierarchical taxonomy that reflects the natural progression from abstract reasoning to embodied action. Our framework organizes physics-aware AI capabilities into four interconnected domains that mirror human cognitive development in physical understanding. Physical Perception represents the ability to intuitively extract physical properties from sensory data through increasingly sophisticated layers, including object recognition, spatial relationships and intrinsic properties. In the inverse direction, Physics Reasoning encompasses symbolic manipulation and theoretical problem-solving, where AI systems leverage mathematical representations to solve physics problems ranging from textbook exercises to research-level challenges. These two complementary pathways converge in World Modeling, where AI systems integrate symbolic knowledge with perceptual understanding to construct predictive models of physical environments, enabling capabilities from video generation to reconstruction. Finally, Embodied Interaction grounds these cognitive capabilities in real-world action, where robots, autonomous vehicles, and navigation agents must reconcile theoretical understanding with the irreversible consequences of physical manipulation. Figure 3 shows our taxonomy. This taxonomy not only delineates current research boundaries but also reveals critical gaps where symbolic reasoning and embodied experience remain disconnected, highlighting the path toward genuinely physics-aware artificial intelligence.
C Physical Perception: From Sensory Data to Physical Understanding
Understanding the physical world begins with perception, where AI systems transform raw sensory inputs into structured representations of physical properties, spatial relationships, dynamic behaviors. Physical perception serves as the essential grounding for subsequent reasoning, modeling, embodied interaction, establishing the observational foundation upon which symbolic reasoning and predictive models are built. While traditional computer vision focuses on semantic recognition, physical perception demands extracting geometry, material properties, force dynamics, causal structure from visual or multimodal observations. This capability enables systems to infer latent physical states that are not directly observable but fundamentally govern real-world phenomena. We organize existing investigations along a progression of increasing cognitive complexity: Object Recognition extracts geometric structure with semantic information, Spatial Perception captures relationships within scene composition, Intrinsic Property Estimation infers material characteristics, Dynamic Estimation predicts motion through temporal evolution, Causal and Counterfactual Inference reveals underlying physical mechanisms. This hierarchical organization reflects how perceptual capabilities build cumulatively, with each level providing foundations for the next, ultimately enabling the transition from observation to reasoning about physical laws.